Please help me solve these two equations:
x+3-2
x+x+2-1x

Answer by stanbon(48546)

x+3-2
x+x+2-1x
What you have posted are not equations.
You cannot "solve" them.
You can,however, simplify them as follows:
x+3-2 = x+1
x+x+2-1x = x+2
cheers,
stan H.